OpenStudy (anonymous):

what is the equation of the function y=5/x translated four units to the left and three units up

OpenStudy (mathmale):

this problem pertains to horizontal and vertical translation. If you translate the graph of the function y=5/x three units to the right, the function becomes \[y=\frac{ 5 }{ x-3 }\]... and if two units to the left, \[y=\frac{ 5 }{ x-(-2) }=\frac{ 5 }{ x+2 }\]

OpenStudy (mathmale):

Please use the info from these 2 examples to write your new formula if the graph of y=5/x is translated four units to the left. Then, think about how you'd change the formula again to represent a vertical shift of +3 units.

OpenStudy (anonymous):

Okay. I was confused as to where in the equation the horisontal translation and the vertical translation go.

OpenStudy (mathmale):

Now you have two examples of horiz. translation. Can you now modify the equation to reflect a horiz. translation of four units to the left?

OpenStudy (anonymous):

y=5/x+4

OpenStudy (mathmale):

\[y=\frac{ 5 }{ x+4 } \]would be better in that there's no doubt that the divisor is (x+4), but yes, you have the concept correct!

OpenStudy (mathmale):

Next, you have to move the whole graph up 3 units. How will the equation change?

OpenStudy (anonymous):

y=(5/(x+4))+3

OpenStudy (mathmale):

Too cool for words!! Ciaran is awesome!

OpenStudy (anonymous):

Thank you! Could you help me with another problem, please?

OpenStudy (mathmale):

Note: Were you to move the whole graph downward by 5 units, you'd get \[y=\frac{ 5 }{ x+4 }-5\]

OpenStudy (mathmale):

Go ahead and post your next question in the "Ask a question" box.

OpenStudy (anonymous):

Oh, okay.